The length of a rectangle is 2 meters less than three times the width.
What is the maximum integer value of the length if the perimeter is less than 43 meters?
A. 5 meters
B. 13 meters
C. 15 meters
D. 16 meters
Could someone please help me with this problem and go over step by step with me?
I believe the answer is 5 meters because it is the only solution which the perimeter is less than 43 meters...??? HELP!!!

Answer by amoresroy(333)

The length of a rectangle is 2 meters less than three times the width.
What is the maximum integer value of the length if the perimeter is less than 43 meters?
Let w = the width of rectangle
L = 3w-2
Perimeter = 2(w+L)
43 = 2(w + 3w-2)
43 = 2(4w-2)
43 = 8w - 4
8w = 47
w = 47/8
L = 3(47/8) - 2
L = 15.6 should be 15
The maximum integer value of the length is 15 meters if the perimeter is less than 43 meters.